OpenStudy (anonymous):

the quantities S and T are positive and are related by the equation S= k/T, where k is a constant. if S increases by 50 % , by what percent the T decreases ? explain

OpenStudy (anonymous):

S gets multiplied by 1.5, so T get divided by 1.5

OpenStudy (anonymous):

k so what's next, can u explain

OpenStudy (anonymous):

decrease in percentage of T = (Told-Tnew)/Told where Tnew=T/1.5; Told=T yelds T(1-1/1.5)/T = 1-1/1.5 -> 33%

OpenStudy (anonymous):

thnk u